Circular No. 7607

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ak IN UGC 11188 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on unfiltered images taken with the 0.5-m Tenagra III automated telescope on Apr. 1.5 (mag ∼ 18.4) and 3.5 UT (mag ∼ 18.1). SN 2001ak is located at α = 18^{h}17^{m}31^{s}.44, δ = +18^{o}54′00′′.8 (equinox 2000.0), which is ∼ 24′′ east and 12′′ south of the diffuse nucleus of UGC 11188. An image taken by the Katzman Automatic Imaging Telescope on 2000 Oct. 8.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001al IN ANONYMOUS GALAXY A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise Ob-

servatory Optical Transients Search (cf. IAUC 6917, 7055) their discovery of the first supernova (R ∼ 21.1) from the second phase of this project on unfiltered CCD images taken by H. Mendelson on Mar. 26 with the widefield SITe camera mounted on the Wise Observatory 1-m telescope. SN 2001al is located at α = 15^{h}44^{m}51^{s}.74, δ = +36^{o}07′29′′.0 (equinox 2000.0), which is in the field of the galaxy cluster Abell 2122 (z = 0.065). No apparent host galaxy is detected. SN 2001al was invisible in images obtained on 2000 Aug. 30 (limiting mag R ∼ 22.0). The new object was confirmed on CCD images taken by J. Dann on Mar. 28 with the Tektronics CCD camera mounted on the same telescope. Finding charts can be obtained at

http://wise-obs.tau.ac.il/~avishay/woots/woots2.html.

A. V. Filippenko, University of California, Berkeley; A. J. Barth, Harvard-Smithsonian Center for Astrophysics; and D. C. Leonard, University of Massachusetts at Amherst, report that inspection of a CCD spectrum (range 390–880 nm) obtained on Mar. 29 UT with the Keck-I telescope reveals that SN 2001al is of type Ia, ∼ 1 month past maximum brightness. The estimated redshift, 0.065–0.07, is consistent with that of the cluster Abell 2122.

CORRIGENDA ‘XTE J1859+226’, IAUC 7506, for J1859+226 read XTE J1859+226 ‘Supernova 2001aj in UGC 10243’, IAUC 7606, line 9, for V. Pozzoli,

Bareggio, Italy, reports read V. Pozzoli, Bareggio, Italy; and G. Sala, Azzate, Italy, report

2001 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green